# CAESAR: Enhancing Federated RL in Heterogeneous MDPs through Convergence-Aware Sampling with Screening

### v1

@@ 4.4. Sampling
Given this intuition, the convergence of value functions emerges as a practical heuristic for estimating whether two agents are learning in the same MDP. This insight leads to our convergence-aware sampling scheme, Sampling, wherein the subset $S$ is determined based on probabilities $p_{i1},\dots,p_{iN}$. Each probability $p_{ij}$ quantifies the likelihood of including agent $j$ in $S$ and is dynamically adjusted based on the observed convergence between $Q_{i}$ and $Q_{j}$. At the onset of training, the server initializes an $N\times N$ matrix $p$, where the entry $p_{ij}$ is set as:
@@
$p_{ij}\leftarrow\begin{cases}\min(p_{ij}+\delta,1)&\text{if }\\ &d(Q_{i}^{(t-H)},Q_{j}^{(t-H)})-d(Q_{i}^{(t)},Q_{j}^{(t)})>\xi,\\ \max(p_{ij}-\delta,0)&\text{otherwise.}\end{cases}$ (6)
@@ 4.5. CAESAR
To implement this, the local performance
$g_{k}\approx\mathbb{E}_{M_{k},\pi_{Q_{k}}}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}_{k}(s_{t},a_{t})\right]$
of each agent $k$ is measured prior to the federated update. During the update, for a given agent $i$, the server initially samples a preliminary subset of agents, $S^{\prime}$, in line with the probabilities $p_{i1},\dots,p_{iN}$. It then further refines $S^{\prime}$ by retaining only those agents whose performance, $g_{j}$, exceeds that of agent $i$ $(g_{j}>g_{i})$. The final subset for aggregation is thus defined as:
$S=\{j:j\in S^{\prime}\text{ and }g_{j}>g_{i}\}$ (7)
@@ 5.1. Experimental Settings
Figure 2. The two GridWorld MDPs. The initial state is 00 in both MDPs. In MDP 1 (top), transiting from state $4$ to $5$ generates a reward of $+1$ and transiting from state $-4$ to $-5$ yields a reward of $-1$. In MDP 2 (bottom), the signs of the rewards are flipped.
@@ 5.2. Hypothesis Verification Using GridWorld
Analysis of Q-Value Evolution. To delve into the effectiveness of the screening process within the CAESAR scheme, we track the progression of Q-values throughout the training period. Fig. 7 and Fig. 8 present the evolution of Q-values for agents assigned to $M_{1}$ under the Sampling and CAESAR schemes, respectively. These plots are generated from training sessions that utilized identical agent configurations and the same random seed. Under Sampling, we notice that only two agents are able to approximate the optimal Q-values (indicated by black dotted lines), while the remaining agents stagnate at suboptimal points. In contrast, when employing CAESAR, a uniform and rapid convergence to optimal values is observed for all agents.

### v2

@@ 4.4. Sampling
Given this intuition, the convergence of value functions emerges as a practical heuristic for estimating whether two agents are learning in the same MDP. This insight leads to our convergence-aware sampling scheme, Sampling, wherein the subset $S$ is sampled based on probabilities $p_{i1},\dots,p_{iN}$. Each probability $p_{ij}$ quantifies the likelihood of including agent $j$ in $S$ and is dynamically adjusted based on the observed convergence between $Q_{i}$ and $Q_{j}$. At the onset of training, the server initializes an $N\times N$ matrix $p$, where the entry $p_{ij}$ is set as:
@@
$p_{ij}\leftarrow\begin{cases}\min(p_{ij}+\delta,1)&\text{if }d(Q_{i}^{(t-H)},Q_{j}^{(t-H)})-d(Q_{i}^{(t)},Q_{j}^{(t)})>\xi,\\ \max(p_{ij}-\delta,0)&\text{otherwise.}\end{cases}$
@@ 4.5. CAESAR
To implement this, the local performance of each agent $k$, $g_{k}\approx\mathbb{E}_{M_{k},\pi_{Q_{k}}}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}_{k}(s_{t},a_{t})\right]$, is measured prior to the federated update. During the update, for a given agent $i$, the server initially samples a preliminary subset of agents, $S^{\prime}$, in line with the probabilities $p_{i1},\dots,p_{iN}$. It then further refines $S^{\prime}$ by retaining only those agents whose performance, $g_{j}$, exceeds that of agent $i$ $(g_{j}>g_{i})$. The final subset for aggregation is thus defined as:
$S=\{j:j\in S^{\prime}\text{ and }g_{j}>g_{i}\}$
@@ 5.1. Experimental Settings
Figure 2. Two GridWorld MDPs. Their initial states are 00. In MDP 1 (top), transiting from state $4$ to $5$ generates a reward of $+1$ and transiting from state $-4$ to $-5$ yields a reward of $-1$. In MDP 2 (bottom), the signs of the rewards are flipped.
@@ 5.2. Hypothesis Verification Using GridWorld
Analysis of Q-Value Evolution. To understand the critical role of the screening process in the CAESAR scheme, we track the progression of Q-values throughout the training period. Fig. 7 and Fig. 8 present the evolution of Q-values for agents assigned to $M_{1}$ under the Sampling and CAESAR schemes, respectively. These plots are generated from training sessions with the same random seed and initial agent configurations. Under Sampling, we notice that only two agents are able to approximate the optimal Q-values (indicated by black dotted lines), while the remaining agents stagnate at suboptimal points. In contrast, when employing CAESAR, a uniform and rapid convergence to optimal values is observed for all agents.